A 150 pound person needs an antibiotic injection with a dosage of 10 mg/kg of body mass. The antibiotic is labeled with a concentration of 1 gram/mL. How many mL of antibiotic sould this person receive?
1
Expert's answer
2019-09-06T04:20:34-0400

Weight of person in "Kg=150 \times 0.453592= 68.04 kg"

Amount of antibiotic injection required"=W \times dosage =68.04 \times 10=680.4 \space mg"

Volume of antibiotic required=680.4/1000"=0.6804 ml"